OpenStudy (anonymous):

1/R=1/R1+1/R2+1/R3 solve for R2, I know its R2 = 1/[1/R - 1/R1 - 1/R3], but how do you get everything divided by 1 in the end?

OpenStudy (anonymous):

is it question of parallel resistance ???

OpenStudy (anonymous):

Yeah, but I'm just solving for R2. Physics practice kinda thing

OpenStudy (anonymous):

if we do the reciprocal r = r1 + r2 + r3

OpenStudy (anonymous):

@li_za12 if we do reciprocal we will have the answer

OpenStudy (anonymous):

R1*R2*R3 --------------- R1*R3 +R1*R2 +R2*R3

OpenStudy (anonymous):

fixed

OpenStudy (anonymous):

but what you do to 1/R2 = 1/R - 1/R1 - 1/R3 to get R2 = 1/ (1/R-1/R1-1/R3)

OpenStudy (phoenixfire):

\[{1 \over R}={1 \over R1}+{1 \over R2}+{1 \over R3}\]Rearrange\[{1 \over R2}={1 \over R}-{1 \over R1}-{1 \over R3}\] Multiply by R2 \[1=R2({1 \over R}-{1 \over R1}-{1 \over R3})\]Divide by 1/R1+1/R3-1/R \[R2={1 \over {{1 \over R}-{1 \over R1}-{1 \over R3}}}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Ah sorry, didn't get what you were asking. I thought you were just asking generally about how to shorthand for Req

OpenStudy (anonymous):

"how do you get everything divided by 1 in the end?" there's a key on your calc. that does that. "1/x" ie "R" "1/x" "+" "-R1" "1/x" "=" "+"-R3" "1/x" "=" "1/x"

OpenStudy (phoenixfire):

@Algebraic! That's not going to solve the equation for R2 for you. All that will do is give you the reciprocal. You still need to do some basic algebra and multiply and divide to isolate R2 on one side of the equation.

OpenStudy (anonymous):

@PhoenixFire It will actually, it's directly from your expression, which is correct.